OpenStudy (calculusxy):

Rachel buys a bag that is on sale for 60% off. If the sale price is $79.98, what is the original price? Show all your work.

OpenStudy (alexandervonhumboldt2):

x*(1-60/100)=79.98

OpenStudy (alexandervonhumboldt2):

79.98/0.4

OpenStudy (alexandervonhumboldt2):

199.95

OpenStudy (calculusxy):

@AlexandervonHumboldt2 Can you explain your equation please?

OpenStudy (alexandervonhumboldt2):

so you decrease the original price by 60%. the original price is not known so x

OpenStudy (alexandervonhumboldt2):

x-x*60/100=x*(1-60/100)

OpenStudy (alexandervonhumboldt2):

x*(1-60/100)=79.98

OpenStudy (calculusxy):

But what about 1-60/100? Why the 1?